# Hyperspherical Latents Improve Continuous-Token Autoregressive Generation

### v1

@@ 1 Introduction
Autoregressive (AR) models have achieved remarkable success in text (radford2018improving, ; brown2020language, ) and have been extended to images (esser2021taming, ; yu2021vector, ), speech (meng2024autoregressive, ), video (teng2025magi, ), and other modalities (lu2025uni, ). Early multimodal AR systems discretized latents with vector quantization (VQ) (gray1984vector, ; van2017neural, ); more recently, continuous-token AR dispenses with codebooks: a VAE (kingma2013auto, ) emits token-level latents and the AR model predicts the next latent in continuous space (e.g., Gaussian mixtures (tschannen2024givt, ) or diffusion objectives (li2024autoregressive, ; sun2024multimodal, )). Yet, when built on the same VAE latents, continuous-token AR models often trail latent diffusion and masked-generation models. [^1] In this paper, “autoregressive” denotes token-by-token generation with unidirectional (causal) self-attention, excluding bidirectional masked/next-scale methods such as MaskGIT (chang2022maskgit, ), MAR (li2024autoregressive, ), and VAR (tian2024visual, ). With the same VAE latents, (li2024autoregressive, ) reports an AR model at FID 4.69 vs. 1.98 for MAR and 2.27 for DiT (peebles2023scalable, ). Prior analyses attribute this gap to variance pathologies during AR decoding (sun2024multimodal, ; team2025nextstep, ): latent variances are heterogeneous across dimensions/tokens and are amplified due to exposure bias and classifier-free guidance (CFG) (ho2022classifier, ), causing stepwise variance drift and collapse. Strengthening the KL term (tschannen2024givt, ) or fixing a large variance (sun2024multimodal, ) improves stability but leaves the root cause intact: scale heterogeneity remains and can still drift during AR decoding with CFG.
We address this with a more principled solution: make all AR inputs and outputs scale-invariant. As illustrated in Fig. 2, the proposed SphereAR couples a hyperspherical VAE (S-VAE) (davidson2018hyperspherical, ; de2020power, ) with an autoregressive Transformer (vaswani2017attention, ) and a token-level diffusion head (li2024autoregressive, ). The S-VAE constrains each latent token to a fixed-radius hypersphere (constant $\ell_{2}$ norm), parameterizing only direction via a unit mean direction vector $\bm{\mu}$ and a scalar concentration $\kappa$. During training, the AR model consumes these hyperspherical latents under teacher forcing. During inference, AR model’s predictions—including those after CFG rescaling—are projected back onto the fixed-radius hypersphere to remove the radial (scale) component. Thus, every signal provided to or produced by the AR model is $\ell_{2}$-normalized to the same radius. A concise theoretical justification supports these design choices, showing why scale-invariant inputs/outputs stabilize AR decoding and why a hyperspherical posterior is preferable to Gaussian alternatives.
@@ 2 Related Work
A large body of work improves the performance of image tokenizers by enhancing reconstruction fidelity and semantic alignment. Typical ingredients include (i) refined training objectives (esser2021taming, ; yao2025reconstruction, ; yang2025latent, ), (ii) CLIP-aligned distillation for better text guidance (peng2022beit, ; qu2025tokenflow, ), and (iii) various decoder improvements (chen2025diffusion, ; yang2025latent, ). These techniques are orthogonal to our approach and can be combined with it. In parallel, a complementary line of work targets the quantization mechanism itself—improving codebook utilization, training stability, and the rate-distortion tradeoff. Building on VQ-VAE (van2017neural, ), extensions include hierarchical VQ-VAE-2 (razavi2019generating, ), residual/hierarchical quantization (lee2022autoregressive, ), and multi-codebook schemes (li2024imagefolder, ). Some methods also adopt spherical or normalized feature geometry in the quantizer: for instance, ViT-VQGAN (yu2021vector, ) normalizes latent features before computing codebook distances, and BSQ (zhao2024image, ) constructs binarized spherical latents for bit-efficient quantization.
By contrast, comparatively less work targets continuous image tokenizers tailored to autoregressive modeling. Most prior approaches follow latent diffusion practice (peebles2023scalable, ) and employ diagonal-Gaussian VAEs. GIVT (tschannen2024givt, ) and LatentLM (sun2024multimodal, ) mitigate instability by inflating or fixing latent variance (e.g., $\beta$-VAE, $\sigma$-VAE), which helps but does not remove scale degrees of freedom. NextStep-1 (team2025nextstep, ) instead normalizes Gaussian-posterior latents to a constant norm, achieving scale invariance. However, both our theoretical analysis and empirical results indicate that hyperspherical posteriors are preferable to post-hoc normalization of diagonal-Gaussian latents.
@@
Autoregressive image generation can be grouped into three families: next-scale, next-set, and next-token prediction. In next-scale prediction (e.g., VAR (tian2024visual, )), images are generated coarse-to-fine across resolutions; within each scale, context is modeled bidirectionally. In next-set prediction (also called masked generation; e.g., MaskGIT (chang2022maskgit, ), MAR (li2024autoregressive, )), a single scale is used and a set of tokens is updated in parallel under bidirectional attention. In next-token prediction (e.g., VQGAN (esser2021taming, ), LlamaGen (sun2024autoregressive, )), the model follows language-style sequence modeling: one token is predicted at a time with strictly unidirectional (causal) attention. We focus on next-token models because they align naturally with autoregressive language modeling and offer headroom for unified multimodal models.
A wide range of next-token variants has been explored: discrete tokens (esser2021taming, ; yu2021vector, ; sun2024autoregressive, ) vs. continuous tokens (tschannen2024givt, ; sun2024multimodal, ); raster order (sun2024autoregressive, ; tschannen2024givt, ) vs. randomized order (pang2025randar, ; yu2024randomized, ); and more (li2025fractal, ). However, at comparable parameter scales, these models have often trailed next-set and next-scale approaches. A key reason is the variance collapse that emerges during autoregressive decoding (sun2024multimodal, ; team2025nextstep, ). We address this by enforcing scale-invariant latents via hyperspherical VAEs, thereby removing scale degrees of freedom. Empirically, this yields substantial gains for sequential AR decoding, with performance that matches or surpasses state-of-the-art next-set and next-scale methods at comparable model budgets.
@@ 3 Method
We observe that with discrete tokens, next-token autoregressive (AR) models can outperform bidirectional masked-generation (MG) approaches. For example, LlamaGen-L (343M, FID 3.07; (sun2024autoregressive, )) vs. MaskGIT (207M, FID 4.02; (chang2022maskgit, )). A system-level study (fan2024fluid, ) further reports that, with discrete tokens, AR consistently achieves better FID than MG across model sizes from 166M to 3.1B. By contrast, with continuous tokens, MG is consistently stronger than AR. This divergence—discrete tokens thriving under AR while continuous tokens do not—motivates us to probe what truly differentiates the two. As illustrated in Fig.3, discrete tokens (Fig.3 a) are normalized on the probability simplex (components sum to 1), yielding scale-invariant inputs and outputs that stabilize AR decoding. In contrast, diagonal-Gaussian latents (Fig.3 b) are unconstrained and can destabilize multi-step AR decoding due to scale drift that compounds across steps. We hypothesize this scale sensitivity is the key issue, and therefore constrain continuous latents to a fixed-radius hypersphere to enforce a constant norm (Fig.3 c). This idea underpins SphereAR: a hyperspherical VAE paired with a causal Transformer equipped with a token-level diffusion head; we detail these components below.
@@ 3.1 From VAE to Hyperspherical VAE
A variational autoencoder (VAE) (kingma2013auto, ) is widely used to compress raw data into a lower-dimensional latent vector. It consists of an encoder $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ that parameterizes an approximate posterior over $\mathbf{z}$ and a decoder $p_{\psi}(\mathbf{x}\mid\mathbf{z})$ that reconstructs $\mathbf{x}$ from $\mathbf{z}$. We train the model by maximizing the evidence lower bound (ELBO):
@@
With this diagonal-Gaussian posterior, the encoder’s scale $\bm{\sigma}_{\phi}(\mathbf{x})$ is data-dependent and per-dimension, yielding heterogeneous variances across dimensions and tokens. This imbalance amplifies exposure bias and can trigger variance collapse in AR decoding, particularly under CFG (sun2024multimodal, ; team2025nextstep, ).
@@
To fully address this issue, we remove the scale degree of freedom in the latent representation, rendering the AR model’s inputs and outputs scale-invariant. Specifically, leveraging hyperspherical VAEs (S-VAEs) (davidson2018hyperspherical, ; de2020power, ), we constrain each latent token to lie on a fixed-radius hypersphere.
For each token, the S-VAE encoder parameterizes a directional posterior on the unit sphere by outputting a unit mean direction $\bm{\mu}_{\phi}(\mathbf{x})\in\mathbb{S}^{d-1}$ (via $\ell_{2}$ normalization; $d$ is the latent dimension) and a nonnegative concentration $\kappa_{\phi}(\mathbf{x})\in\mathbb{R}_{\geq 0}$. For notational convenience, let $\bm{\mu}=\bm{\mu}_{\phi}(\mathbf{x})$ and $\kappa=\kappa_{\phi}(\mathbf{x})$. S-VAE adopts a von Mises–Fisher (vMF) distribution (davidson2018hyperspherical, ) for the directional approximate posterior:
@@
While vMF is principled for spherical latents, it can be less efficient due to the need for rejection sampling. For efficiency, we adopt the Power Spherical posterior (de2020power, ) on $\mathbb{S}^{d-1}$,
@@ 3.2 Continuous-Token Autoregressive Transformer
To predict the next continuous token $\mathbf{z}_{k}$, we follow MAR (li2024autoregressive, ) and attach a token-level diffusion head. Conditioned on $\mathbf{h}_{k-1}$, the head progressively transforms a simple prior (e.g., $\mathcal{N}(\mathbf{0},\mathbf{I})$) into the data distribution of the next token $\mathbf{z}_{k}$.
We train the diffusion head with Rectified Flow (lipman2022flow, ; liu2022flow, ). Given a prior $\mathbf{z}_{k}^{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, target $\mathbf{z}_{k}^{1}=\mathbf{z}_{k}$, and a continuous time $t\in(0,1)$, we form the linear interpolation
@@ 3.3 Model Architectures
Although VQGAN‐style CNN backbones (esser2021taming, ) are effective for latent VAEs, their throughput is limited by large convolutional activation maps. To improve efficiency without sacrificing quality, we adopt a hybrid backbone: the encoder uses a lightweight CNN stem with downsampling for patchification, followed by a stack of Transformer blocks; the decoder mirrors this with a Transformer stack that refines latent tokens and a lightweight CNN with upsampling for unpatchification and pixel reconstruction. This preserves the CNN’s strong local inductive bias while leveraging the Transformer’s efficient global modeling at token resolution, yielding a favorable speed–quality trade-off. As shown in Appendix D, the hybrid matches CNN baselines in quality while being about $2.6\times$ faster.
@@
Following prior work (sun2024autoregressive, ; sun2024multimodal, ), we adopt a modern causal Transformer. Concretely, we use pre-norm Transformer blocks with RMSNorm (zhang2019root, ; xiong2020layer, ), FlashAttention for efficient attention computation (dao2022flashattention, ), and SwiGLU feed-forward layers (shazeer2020glu, ). For image positional encoding, we employ 2D rotary embeddings (RoPE) (su2024roformer, ) applied in raster-scan order. All self-attention is strictly unidirectional (causal mask). For the diffusion head, we follow MAR (li2024autoregressive, ) and use an MLP architecture.
@@ 4 Experiments
We evaluate SphereAR on ImageNet-1K (deng2009imagenet, ) class-conditional generation of a resolution of $\times256$, comparing against previous strong baselines. Beyond end-to-end comparisons, we include targeted studies to substantiate our design choices, focusing on the following questions: (1) S-VAE vs. diagonal-Gaussian: Does S-VAE outperform diagonal-Gaussian VAEs for continuous-token AR? (2) Post-hoc normalization: If we apply $\ell_{2}$ normalization to latents from a diagonal-Gaussian VAE, how does it compare with S-VAE? (3) Component contributions: Which parts of S-VAE drive the gains—(i) the hyperspherical posterior, (ii) normalization applied to the VAE decoder input, or (iii) normalization applied to AR inputs/outputs?
@@ 4.1 Implementation Details
We adopt a Power Spherical de2020power directional posterior with latent dimensionality $d=16$ and fix the radius to $R=\sqrt{d}$. Complete setting for the S-VAE’s backbone is provided in Appendix D. We train S-VAE from scratch on ImageNet-1K (deng2009imagenet, ) with random-crop augmentation, optimizing a weighted sum of ELBO (reconstruction ++ KL), perceptual (johnson2016perceptual, ; zhang2018unreasonable, ), and adversarial (isola2017image, ) losses. Optimization uses AdamW (kingma2014adam, ; loshchilov2017decoupled, ) for 100 epochs (batch size $256$, learning rate $1\times 10^{-4}$, $\beta=(0.9,0.95)$, weight decay $0.05$).
@@
Following MAR (li2024autoregressive, ), we instantiate three model sizes for SphereAR. SphereAR-B uses 24 Transformer blocks (hidden size 768) and a diffusion head with 6 feed-forward blocks (hidden size 768). SphereAR-L uses 32 Transformer blocks (hidden size 1024) and a diffusion head with 8 feed-forward blocks (hidden size 1024). SphereAR-H uses 40 Transformer blocks (hidden size 1280) and a diffusion head with 12 feed-forward blocks (hidden size 1280). As in MAR, we employ multiple class-conditioning tokens (16 in our models vs. 64 in MAR) and apply class-token dropout with probability $0.1$ during training to enable classifier-free guidance (CFG) at inference. Models are trained on ImageNet-1K for 400 epochs using S-VAE latents with AdamW (batch size $512$, $\beta=(0.9,0.95)$, weight decay $0.05$), a cosine learning-rate schedule with 20k linear warmup steps and peak learning rate $3\times 10^{-4}$, and an exponential moving average (EMA) of weights with decay $0.9999$. Under these settings, SphereAR-B, SphereAR-L and SphereAR-H contain $\sim$208M, $\sim$479M and $\sim$943M parameters, respectively.

### v2

@@ 1 Introduction
Autoregressive (AR) models have achieved remarkable success in text (Radford et al., 2018; Brown et al., 2020) and have been extended to images (Esser et al., 2021; Yu et al., 2021), speech (Meng et al., 2024), video (Teng et al., 2025), and other modalities (Lu et al., 2025). Early multimodal AR systems discretized latents with vector quantization (VQ) (Gray, 1984; Van Den Oord et al., 2017); more recently, continuous-token AR dispenses with codebooks: a VAE (Kingma and Welling, 2013) emits token-level latents and the AR model predicts the next latent in continuous space (e.g., Gaussian mixtures (Tschannen et al., 2024) or diffusion objectives (Li et al., 2024a; Sun et al., 2024b)). Yet, when built on the same VAE latents, continuous-token AR models often trail latent diffusion and masked-generation models. [^1] In this paper, “autoregressive” denotes token-by-token generation with unidirectional (causal) self-attention, excluding bidirectional masked/next-scale methods such as MaskGIT (Chang et al., 2022), MAR (Li et al., 2024a), and VAR (Tian et al., 2024). With the same VAE latents, Li et al. (2024a) report an AR model at FID 4.69 vs. 1.98 for MAR and 2.27 for DiT (Peebles and Xie, 2023). Prior analyses attribute this gap to variance pathologies during AR decoding (Sun et al., 2024b; Team et al., 2025): latent variances are heterogeneous across dimensions/tokens and are amplified due to exposure bias and classifier-free guidance (CFG) (Ho and Salimans, 2022), causing stepwise variance drift and collapse. Strengthening the KL term (Tschannen et al., 2024) or fixing a large variance (Sun et al., 2024b) improves stability but leaves the root cause intact: scale heterogeneity remains and can still drift during AR decoding with CFG.
We address this with a more principled solution: make all AR inputs and outputs scale-invariant. As illustrated in Fig. 2, the proposed SphereAR couples a hyperspherical VAE (S-VAE) (Davidson et al., 2018; De Cao and Aziz, 2020) with an autoregressive Transformer (Vaswani et al., 2017) and a token-level diffusion head (Li et al., 2024a). The S-VAE constrains each latent token to a fixed-radius hypersphere (constant $\ell_{2}$ norm), parameterizing only direction via a unit mean direction vector $\bm{\mu}$ and a scalar concentration $\kappa$. During training, the AR model consumes these hyperspherical latents under teacher forcing. During inference, AR model’s predictions—including those after CFG rescaling—are projected back onto the fixed-radius hypersphere to remove the radial (scale) component. Thus, every signal provided to or produced by the AR model is $\ell_{2}$-normalized to the same radius. A concise theoretical justification supports these design choices, showing why scale-invariant inputs/outputs stabilize AR decoding and why a hyperspherical posterior is preferable to Gaussian alternatives.
@@ 2 Related Work
A large body of work improves the performance of image tokenizers by enhancing reconstruction fidelity and semantic alignment. Typical ingredients include (i) refined training objectives (Esser et al., 2021; Yao et al., 2025; Yang et al., 2025), (ii) CLIP-aligned distillation for better text guidance (Peng et al., 2022; Qu et al., 2025), and (iii) various decoder improvements (Chen et al., 2025; Yang et al., 2025). These techniques are orthogonal to our approach and can be combined with it. In parallel, a complementary line of work targets the quantization mechanism itself—improving codebook utilization, training stability, and the rate-distortion tradeoff. Building on VQ-VAE (Van Den Oord et al., 2017), extensions include hierarchical VQ-VAE-2 (Razavi et al., 2019), residual/hierarchical quantization (Lee et al., 2022), and multi-codebook schemes (Li et al., 2024b). Some methods also adopt spherical or normalized feature geometry in the quantizer: for instance, ViT-VQGAN (Yu et al., 2021) normalizes latent features before computing codebook distances, and BSQ (Zhao et al., 2024) constructs binarized spherical latents for bit-efficient quantization.
By contrast, comparatively less work targets continuous image tokenizers tailored to autoregressive modeling. Most prior approaches follow latent diffusion practice (Peebles and Xie, 2023) and employ diagonal-Gaussian VAEs. GIVT (Tschannen et al., 2024) and LatentLM (Sun et al., 2024b) mitigate instability by inflating or fixing latent variance (e.g., $\beta$-VAE, $\sigma$-VAE), which helps but does not remove scale degrees of freedom. NextStep-1 (Team et al., 2025) instead normalizes Gaussian-posterior latents to a constant norm, achieving scale invariance. However, both our theoretical analysis and empirical results indicate that hyperspherical posteriors are preferable to post-hoc normalization of diagonal-Gaussian latents.
@@
Autoregressive image generation can be grouped into three families: next-scale, next-set, and next-token prediction. In next-scale prediction (e.g., VAR (Tian et al., 2024)), images are generated coarse-to-fine across resolutions; within each scale, context is modeled bidirectionally. In next-set prediction (also called masked generation; e.g., MaskGIT (Chang et al., 2022), MAR (Li et al., 2024a)), a single scale is used and a set of tokens is updated in parallel under bidirectional attention. In next-token prediction (e.g., VQGAN (Esser et al., 2021), LlamaGen (Sun et al., 2024a)), the model follows language-style sequence modeling: one token is predicted at a time with strictly unidirectional (causal) attention. We focus on next-token models because they align naturally with autoregressive language modeling and offer headroom for unified multimodal models.
A wide range of next-token variants has been explored: discrete tokens (Esser et al., 2021; Yu et al., 2021; Sun et al., 2024a) vs. continuous tokens (Tschannen et al., 2024; Sun et al., 2024b); raster order (Sun et al., 2024a; Tschannen et al., 2024) vs. randomized order (Pang et al., 2025; Yu et al., 2024); and more (Li et al., 2025). However, at comparable parameter scales, these models have often trailed next-set and next-scale approaches. A key reason is the variance collapse that emerges during autoregressive decoding (Sun et al., 2024b; Team et al., 2025). We address this by enforcing scale-invariant latents via hyperspherical VAEs, thereby removing scale degrees of freedom. Empirically, this yields substantial gains for sequential AR decoding, with performance that matches or surpasses state-of-the-art next-set and next-scale methods at comparable model budgets.
@@ 3 Method
We observe that with discrete tokens, next-token autoregressive (AR) models can outperform bidirectional masked-generation (MG) approaches. For example, LlamaGen-L (343M, FID 3.07; (Sun et al., 2024a)) vs. MaskGIT (207M, FID 4.02; (Chang et al., 2022)). A system-level study (Fan et al., 2024) further reports that, with discrete tokens, AR consistently achieves better FID than MG across model sizes from 166M to 3.1B. By contrast, with continuous tokens, MG is consistently stronger than AR. This divergence—discrete tokens thriving under AR while continuous tokens do not—motivates us to probe what truly differentiates the two. As illustrated in Fig.3, discrete tokens (Fig.3 a) are normalized on the probability simplex (components sum to 1), yielding scale-invariant inputs and outputs that stabilize AR decoding. In contrast, diagonal-Gaussian latents (Fig.3 b) are unconstrained and can destabilize multi-step AR decoding due to scale drift that compounds across steps. We hypothesize this scale sensitivity is the key issue, and therefore constrain continuous latents to a fixed-radius hypersphere to enforce a constant norm (Fig.3 c). This idea underpins SphereAR: a hyperspherical VAE paired with a causal Transformer equipped with a token-level diffusion head; we detail these components below.
@@ 3.1 From VAE to Hyperspherical VAE
A variational autoencoder (VAE) (Kingma and Welling, 2013) is widely used to compress raw data into a lower-dimensional latent vector. It consists of an encoder $q_{\phi}(\mathbf{z}\mid\mathbf{x})$ that parameterizes an approximate posterior over $\mathbf{z}$ and a decoder $p_{\psi}(\mathbf{x}\mid\mathbf{z})$ that reconstructs $\mathbf{x}$ from $\mathbf{z}$. We train the model by maximizing the evidence lower bound (ELBO):
@@
With this diagonal-Gaussian posterior, the encoder’s scale $\bm{\sigma}_{\phi}(\mathbf{x})$ is data-dependent and per-dimension, yielding heterogeneous variances across dimensions and tokens. This imbalance amplifies exposure bias and can trigger variance collapse in AR decoding, particularly under CFG (Sun et al., 2024b; Team et al., 2025).
@@
To fully address this issue, we remove the scale degree of freedom in the latent representation, rendering the AR model’s inputs and outputs scale-invariant. Specifically, leveraging hyperspherical VAEs (S-VAEs) (Davidson et al., 2018; De Cao and Aziz, 2020), we constrain each latent token to lie on a fixed-radius hypersphere.
For each token, the S-VAE encoder parameterizes a directional posterior on the unit sphere by outputting a unit mean direction $\bm{\mu}_{\phi}(\mathbf{x})\in\mathbb{S}^{d-1}$ (via $\ell_{2}$ normalization; $d$ is the latent dimension) and a nonnegative concentration $\kappa_{\phi}(\mathbf{x})\in\mathbb{R}_{\geq 0}$. For notational convenience, let $\bm{\mu}=\bm{\mu}_{\phi}(\mathbf{x})$ and $\kappa=\kappa_{\phi}(\mathbf{x})$. S-VAE adopts a von Mises–Fisher (vMF) distribution (Davidson et al., 2018) for the directional approximate posterior:
@@
While vMF is principled for spherical latents, it can be less efficient due to the need for rejection sampling. For efficiency, we adopt the Power Spherical posterior (De Cao and Aziz, 2020) on $\mathbb{S}^{d-1}$,
@@ 3.2 Continuous-Token Autoregressive Transformer
To predict the next continuous token $\mathbf{z}_{k}$, we follow MAR (Li et al., 2024a) and attach a token-level diffusion head. Conditioned on $\mathbf{h}_{k-1}$, the head progressively transforms a simple prior (e.g., $\mathcal{N}(\mathbf{0},\mathbf{I})$) into the data distribution of the next token $\mathbf{z}_{k}$.
We train the diffusion head with Rectified Flow (Lipman et al., 2022; Liu et al., 2022). Given a prior $\mathbf{z}_{k}^{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, target $\mathbf{z}_{k}^{1}=\mathbf{z}_{k}$, and a continuous time $t\in(0,1)$, we form the linear interpolation
@@ 3.3 Model Architectures
Although VQGAN‐style CNN backbones (Esser et al., 2021) are effective for latent VAEs, their throughput is limited by large convolutional activation maps. To improve efficiency without sacrificing quality, we adopt a hybrid backbone: the encoder uses a lightweight CNN stem with downsampling for patchification, followed by a stack of Transformer blocks; the decoder mirrors this with a Transformer stack that refines latent tokens and a lightweight CNN with upsampling for unpatchification and pixel reconstruction. This preserves the CNN’s strong local inductive bias while leveraging the Transformer’s efficient global modeling at token resolution, yielding a favorable speed–quality trade-off. As shown in Appendix D, the hybrid matches CNN baselines in quality while being about $2.6\times$ faster.
@@
Following prior work (Sun et al., 2024a; b), we adopt a modern causal Transformer. Concretely, we use pre-norm Transformer blocks with RMSNorm (Zhang and Sennrich, 2019; Xiong et al., 2020), FlashAttention for efficient attention computation (Dao et al., 2022), and SwiGLU feed-forward layers (Shazeer, 2020). For image positional encoding, we employ 2D rotary embeddings (RoPE) (Su et al., 2024) applied in raster-scan order. All self-attention is strictly unidirectional (causal mask). For the diffusion head, we follow MAR (Li et al., 2024a) and use an MLP architecture.
@@ 4 Experiments
We evaluate SphereAR on ImageNet-1K (Deng et al., 2009) class-conditional generation of a resolution of $\times256$, comparing against previous strong baselines. Beyond end-to-end comparisons, we include targeted studies to substantiate our design choices, focusing on the following questions: (1) S-VAE vs. diagonal-Gaussian: Does S-VAE outperform diagonal-Gaussian VAEs for continuous-token AR? (2) Post-hoc normalization: If we apply $\ell_{2}$ normalization to latents from a diagonal-Gaussian VAE, how does it compare with S-VAE? (3) Component contributions: Which parts of S-VAE drive the gains—(i) the hyperspherical posterior, (ii) normalization applied to the VAE decoder input, or (iii) normalization applied to AR inputs/outputs?
@@ 4.1 Implementation Details
We adopt a Power Spherical De Cao and Aziz (2020) directional posterior with latent dimensionality $d=16$ and fix the radius to $R=\sqrt{d}$. Complete setting for the S-VAE’s backbone is provided in Appendix D. We train S-VAE from scratch on ImageNet-1K (Deng et al., 2009) with random-crop augmentation, optimizing a weighted sum of ELBO (reconstruction ++ KL), perceptual (Johnson et al., 2016; Zhang et al., 2018), and adversarial (Isola et al., 2017) losses. Optimization uses AdamW (Kingma, 2014; Loshchilov and Hutter, 2017) for 100 epochs (batch size $256$, learning rate $1\times 10^{-4}$, $\beta=(0.9,0.95)$, weight decay $0.05$).
@@
Following MAR (Li et al., 2024a), we instantiate three model sizes for SphereAR. SphereAR-B uses 24 Transformer blocks (hidden size 768) and a diffusion head with 6 feed-forward blocks (hidden size 768). SphereAR-L uses 32 Transformer blocks (hidden size 1024) and a diffusion head with 8 feed-forward blocks (hidden size 1024). SphereAR-H uses 40 Transformer blocks (hidden size 1280) and a diffusion head with 12 feed-forward blocks (hidden size 1280). As in MAR, we employ multiple class-conditioning tokens (16 in our models vs. 64 in MAR) and apply class-token dropout with probability $0.1$ during training to enable classifier-free guidance (CFG) at inference. Models are trained on ImageNet-1K for 400 epochs using S-VAE latents with AdamW (batch size $512$, $\beta=(0.9,0.95)$, weight decay $0.05$), a cosine learning-rate schedule with 20k linear warmup steps and peak learning rate $3\times 10^{-4}$, and an exponential moving average (EMA) of weights with decay $0.9999$. Under these settings, SphereAR-B, SphereAR-L and SphereAR-H contain $\sim$208M, $\sim$479M and $\sim$943M parameters, respectively.